OpenStudy (anonymous):

Factor x^2 - 25. A. (x + 5)(x - 5) B. (x - 5)(x - 5) C. (x + 5)(x + 5) D. None of the above

jimthompson5910 (jim_thompson5910):

hints: 25 = 5^2 a^2 - b^2 = (a-b)(a+b)

OpenStudy (anonymous):

C.

OpenStudy (anonymous):

D. sorry I read it wrong.

jimthompson5910 (jim_thompson5910):

another hint x^2 - 25 is the same as x^2 - 5^2

jimthompson5910 (jim_thompson5910):

use the rule a^2 - b^2 = (a-b)(a+b)

OpenStudy (isaiah.feynman):

@AntoinetteWillman Its A!